Susie cycled to a park at an average speed of 10km/h. her uncle drove her home from the park at an average speed of 50km/h. her trip back took 1.6h less than herr trip to the park. Find the distance between Susie's home and the park.
Answer by checkley77(12844)

d=rt or rt=rt seeing as the d is the same.
10x=50(x-1.6)
10x=50x-80
10x-50x=-80
-40x=-80
x=2 hours is the time for biking the park.
10*2=20 km is the distance between home & the park.
proof:
50(2-1.6)=50*.4=20 km is the distance of the ride home.